Q.3: Blue Beverages Inc. has bonds outstanding of Rs 5,000 par value (Face value or maturity value), which are currently selling at par (face value). The bonds have 5 years to maturity and a 10% coupon rate with interest paid semiannually. Green Beverages Inc. has a very similar bond issue outstanding. Every bond feature of Green Beverages is the same as for the Blue Beverages except that Green’s bonds mature in 15 years. The market’s nominal annual required rate of return for both bond issues has recently faced a sudden fall to 8%.

Required:

a. Which beverage’s bonds would show the greatest price change and why? Support your answer with conceptual rationale. (Theoretical but logical explanation is required in this part of the question)

b. Determine the bond price for each beverage company’s bonds at the market new required rate of return. Which bond’s price increased the most and by how much?



1
Expert's answer
2021-07-01T05:04:22-0400

a) Green Beverages bonds will show a larger change in price. This is associated with a longer maturation period, which will lead to a larger discounting.

b) "Price_{Blue}=500*\frac{\frac{1}{1.08}*(1-\frac{1}{1.08}^5)}{1-\frac{1}{1.08}}+\frac{5000}{1.08^5}=5399.27"

"Price_{Green}=500*\frac{\frac{1}{1.08}*(1-\frac{1}{1.08}^{15})}{1-\frac{1}{1.08}}+\frac{5000}{1.08^{15}}=5855.95"

Green Beverages bonds increased more because have more coupon income ( because of more number of payments).